# Fully Dynamic Euclidean $k$-Means

### v2

@@ 9 Approximate Range Query Structure
$\varphi(\textsc{Ball}(x,r))\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \Phi_{x}\cap\varphi([\Delta]^{d})\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \varphi(\textsc{Ball}(x,2\Gamma\cdot r)).$ (17)
@@ 9.2 Application: Approximate Nearest Neighbor Structure
As a warm-up application of Lemma 9.1, we describe how to design an approximate nearest neighbor (ANN) oracle – a data structure that, given a dynamic center set $S\subseteq[\Delta]^{d}$, returns for each query point $x$ a (distinct) approximate nearest center $s\in S-x$. We note that similar guarantee of ANN has also been obtained in e.g., [BDF+24], via more specific techniques albeit it may potentially yield better tradeoff beyond consistent hashing.
@@
For any $10<\epsilon<$ 1, there exists a data structure that maintains a dynamic center set $S\subseteq[\Delta]^{d}$, and for every query $x\in[\Delta]^{d}$, returns a center $s\in S-x$ such that $\mathrm{dist}(x,s)\leq O(\epsilon^{-3/2})\cdot\mathrm{dist}(x,S-x)$. The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has update and query time $\tilde{O}(2^{\epsilon d})$.
@@
An ANN oracle can be used to derive many useful data structures. By plugging Lemma 9.2 into the framework of [BGJ+24], we obtain the following ANN indicator data structure.
@@
For any $\gamma\geq 0$ and $10<\epsilon<$ 1, there exists a data structure that maintains a dynamic set $S\subseteq[\Delta]^{d}$ and for every $s\in S$, maintains a bit $b_{\gamma}(s,S)\in\{0,1\}$ with the following guarantee.
@@
Follows from Lemma 9.2 and [BGJ+24, Theorem 3.1]
@@
For any $10<\epsilon<$ 1, there exists a data structure that maintains a dynamic set $S\subseteq[\Delta]^{d}$ and for every $s\in S$, maintains a value $\widehat{\mathrm{dist}}(s,S-s)$ such that
@@
First, we show the correctness of $\{\widehat{\mathrm{dist}}(s,S-s)\}_{s\in S}$. Consider a specific center $s\in S$. By construction, we have $b_{\ell_{s}}^{s}=1\implies\mathrm{dist}(s,S-s)\leq\epsilon^{-3/2}\cdot 2^{\ell_{s}}=\widehat{\mathrm{dist}}(s,S-s)$. When $\ell_{s}=1$, we have $\mathrm{dist}(s,S-s)\geq 1=\frac{\widehat{\mathrm{dist}}(s,S-s)}{2\epsilon^{-3/2}}$, since two distinct points in the discrete space $[\Delta]^{d}$ have distance $\geq 1$. When $\ell_{s}\neq 1$, we have $b_{\ell_{s}-1}^{s}=0\implies\mathrm{dist}(s,S-s)>2^{\ell_{s}-1}=\frac{\widehat{\mathrm{dist}}(s,S-s)}{2\epsilon^{-3/2}}$. In either case, we always have $\mathrm{dist}(s,S-s)\leq\widehat{\mathrm{dist}}(s,S-s)\leq 2\epsilon^{-3/2}\cdot\mathrm{dist}(s,S-s)$.
@@ 9.3 Application: $1$-Means Range Query
Here, we use Lemma 9.1 to devise a data structure that supports estimating both the $1$-means objective and its solution on any given approximate ball. This data structure is obtained by plugging a dynamic coreset algorithm from [HK20] (for the entire set $X$) into Lemma 9.1; hence, for each query, we first obtain a small-sized coreset and then compute the objective and/or solution based on it.
@@
For any $10<\epsilon<$ 1, there exists a data structure that maintains a dynamic weighted dataset $X\subseteq[\Delta]^{d}$, and for every query point $x\in[\Delta]^{d}$ and radius $r\geq 0$, returns a real number $b_{x}\geq 0$, a center $c^{*}\in[\Delta]^{d}$, and two estimates $\widehat{\textsc{cost}}_{c^{*}}$ and $\widehat{\textsc{cost}}_{x}$, such that there exists a subset $B_{x}\subseteq X$ satisfying the following conditions.
@@
We first consider two extreme cases of the query radius $r$. In the case of a small radius $10\leq r<$ 1, we have $\textsc{Ball}(x,r)=\{x\}$ (recall that any two distinct points in the discrete space $[\Delta]^{d}$ have distance at least $1$), and we simply consider the point subset $B_{x}=\{x\}\cap X$. Either subcase – $B_{x}=\{x\}$ or $B_{x}=\emptyset$ – is trivial. In the case of a large radius $r>\sqrt{d}\Delta=\mathrm{diam}([\Delta]^{d})$, we have $\textsc{Ball}(x,r)\cap X=X$, and we simply consider the entire point set $B_{x}=X$. Thus, this extreme case $r>\sqrt{d}\Delta$ reduces to the boundary case $r=\sqrt{d}\Delta$.
@@
The following claim provides a data structure that (1) maintains the total weight of the dataset (a trivial task), and (2) constructs a coreset for the entire dataset, following the approach of [HK20].101010[HK20] provides a general $(1+\epsilon)$-coreset construction for $k$-means. In this paper, we only need a constant-factor coreset for $1$-means (e.g., $k=1$, $\epsilon=0.5$), so we state this special case.
@@
  • for every $j\in[t]$, the following holds with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$. [^11] While Lemma 9.1 does not explicitly consider success probability, its analysis can be easily adapted to handle data structures that succeed with high probability.
@@ 10.1 Data Structure Description
In this part, we describe our data structure for maintaining an approximate assignment. We first present static and dynamic objects that are explicitly maintained by our data structure (Definition 10.1). Then, we show how these explicit objects define an implicit partition of the non-center points $X-S$ (Claim 10.5). Built on top of this implicit partition, we define an implicit assignment $\sigma:X-S\to S$ (Claim 10.7).
@@
    $\big{\{}s\in S:\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq\tfrac{\rho_{i}}{\Gamma}\big{\}}\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ S(i,z)\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \big{\{}s\in S:\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq 2\rho_{i}\big{\}}.$ (21)
@@
Initially, both $X$ and $S$ are empty =absent=$\emptyset$= ∅, so the initialization only needs to compute the $m+1$ hash functions $\{\varphi_{i}\}_{i\in[0,m]}$. By Lemma 8.2, every $\varphi_{i}$ can be constructed in $\tilde{O}(1)$ time with a success probability of $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$. Thus, the total initialization time is $(m+1)\cdot\tilde{O}(1)=\tilde{O}(1)$. Moreover, taking a union bound, the overall success probability is $1-\frac{m+1}{\mathrm{poly}(\Delta^{d})}=1-\frac{1}{\mathrm{poly}(\Delta^{d})}$. ∎
@@
Finally, we show how the explicit objects (Definition 10.1) implicitly define a partition of the non-center points $X-S$ (Claim 10.5). This partition exhibits an approximate equidistant structure, where points in each part are roughly equally distant from the center set (Claim 10.6). We further define an implicit approximate assignment based on this partition (Claim 10.7).
@@
$\mathrm{dist}(s,x)$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(s,\varphi_{i-1}^{-1}(z_{i-1}))+\mathrm{diam}(\varphi_{i-1}^{-1}(z_{i-1}))$ (Triangle inequality)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ 2\rho_{i-1}+\rho_{i-1}$ (Item 2 of Definition 10.1)
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \rho_{i}/\Gamma.$ (Item 1 of Definition 10.1)
Then, $\mathrm{dist}(s,\varphi_{i}^{-1}(z_{i}))\leq\mathrm{dist}(s,x)\leq\rho_{i}/\Gamma\implies s\in S(i,z_{i})=S(i,x)$ by Equations 21 and 22. The arbitrariness of $s\in S(i-1,x)$ implies that $S(i-1,x)\subseteq S(i,x)$. This finishes the proof. ∎
@@
$\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ is a partition of $X-S$.
@@
$\frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,S)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,\sigma^{\prime}(i,z))\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
@@
[…]
@@
$\mathrm{dist}(x,S)$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,\sigma^{\prime}(i,z))$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(\sigma^{\prime}(i,z),\varphi_{i}^{-1}(z))+\mathrm{diam}(\varphi_{i}^{-1}(z))$ ($\varphi_{i}(x)=z$)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ (3\Gamma)^{i}+\frac{1}{2}\cdot(3\Gamma)^{i}$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
@@
$\min_{s\in S}\mathrm{dist}(s,\varphi_{i-1}^{-1}(z^{\prime}))\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{\rho_{i}}{\Gamma}\leavevmode\nobreak\ >\leavevmode\nobreak\ \frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}.$
@@
$\mathrm{dist}(x,S)\leavevmode\nobreak\ =\leavevmode\nobreak\ \min_{s\in S}\mathrm{dist}(x,s)\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \min_{s\in S}\mathrm{dist}(s,\varphi_{i-1}^{-1}(z^{\prime}))\leavevmode\nobreak\ >\leavevmode\nobreak\ \frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}.$
@@
$\frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,S)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \mathrm{dist}(x,\sigma^{\prime}(i,z))\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
@@
The following Claim 10.7 provides an implicit scheme for maintaining an approximate assignment, which follows directly from Claims 10.5 and 10.6.
@@ 10.2 Update Time Analysis
$\varphi_{i}(\textsc{Ball}(s,\tfrac{\rho_{i}}{\Gamma}))\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \Phi_{i}\cap\varphi_{i}([\Delta]^{d})\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \varphi_{i}(\textsc{Ball}(s,2\rho_{i})).$
@@
$z\leavevmode\nobreak\ =\leavevmode\nobreak\ \varphi_{i}(y)\leavevmode\nobreak\ \in\leavevmode\nobreak\ \varphi_{i}(\textsc{Ball}(x,(3\Gamma)^{i-1}))\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \varphi_{i}(\textsc{Ball}(x,\tfrac{1}{\Gamma}\cdot(3\Gamma)^{i}))\leavevmode\nobreak\ \subseteq\leavevmode\nobreak\ \Phi_{x}.$
@@
The total update time for the data structure of Definition 10.1 follows by combining the results from Claims 10.9, 10.10, 10.11, 10.12 and 10.13. Among these, the most expensive one is maintaining the mappings $\{f(i,z)\}_{(i,z)\in H}$, which requires $\tilde{O}(\Lambda^{2})$ time. Therefore, the overall update time for the data structure is $\tilde{O}(\Lambda^{2})$. ∎

### v3

@@ 10 Approximate Range Query Structure
$\varphi(\textsc{Ball}(x,r))\penalty 10000\ \subseteq\penalty 10000\ \Phi_{x}\cap\varphi([\Delta]^{d})\penalty 10000\ \subseteq\penalty 10000\ \varphi(\textsc{Ball}(x,2\Gamma\cdot r)).$ (18)
@@ 10.2 Application: Approximate Nearest Neighbor Structure
As a warm-up application of Lemma 10.1, we describe how to design an approximate nearest neighbor (ANN) oracle – a data structure that, given a dynamic center set $S\subseteq[\Delta]^{d}$, returns for each query point $x$ a (distinct) approximate nearest center $s\in S-x$. We note that similar guarantee of ANN has also been obtained in e.g., [DBLP:conf/nips/0001DFGHJL24], via more specific techniques albeit it may potentially yield better tradeoff beyond consistent hashing.
@@
For any $0<\epsilon<1$, there exists a data structure that maintains a dynamic center set $S\subseteq[\Delta]^{d}$, and for every query $x\in[\Delta]^{d}$, returns a center $s\in S-x$ such that $\mathrm{dist}(x,s)\leq O(\epsilon^{-3/2})\cdot\mathrm{dist}(x,S-x)$. The data structure works against an adaptive adversary and, with probability at least $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$, has update and query time $\tilde{O}(2^{\epsilon d})$.
@@
An ANN oracle can be used to derive many useful data structures. By plugging Lemma 10.2 into the framework of [BhattacharyaGJQ24], we obtain the following ANN indicator data structure.
@@
For any $\gamma\geq 0$ and $0<\epsilon<1$, there exists a data structure that maintains a dynamic set $S\subseteq[\Delta]^{d}$ and for every $s\in S$, maintains a bit $b_{\gamma}(s,S)\in\{0,1\}$ with the following guarantee.
@@
Follows from Lemma 10.2 and [BhattacharyaGJQ24, Theorem 3.1]
@@
For any $0<\epsilon<1$, there exists a data structure that maintains a dynamic set $S\subseteq[\Delta]^{d}$ and for every $s\in S$, maintains a value $\widehat{\mathrm{dist}}(s,S-s)$ such that
@@
First, we show the correctness of $\{\widehat{\mathrm{dist}}(s,S-s)\}_{s\in S}$. Consider a specific center $s\in S$. By construction, we have $b_{\ell_{s}}^{s}=1\implies\mathrm{dist}(s,S-s)\leq\epsilon^{-3/2}\cdot 2^{\ell_{s}}=\widehat{\mathrm{dist}}(s,S-s)$. When $\ell_{s}=1$, we have $\mathrm{dist}(s,S-s)\geq 1=\frac{\widehat{\mathrm{dist}}(s,S-s)}{2\epsilon^{-3/2}}$, since two distinct points in the discrete space $\Delta]^{d}$ have distance $\geq 1$. When $\ell_{s}\neq 1$, we have $b_{\ell_{s}-1}^{s}=0\implies\mathrm{dist}(s,S-s)>2^{\ell_{s}-1}=\frac{\widehat{\mathrm{dist}}(s,S-s)}{2\epsilon^{-3/2}}$. In either case, we always have $\mathrm{dist}(s,S-s)\leq\widehat{\mathrm{dist}}(s,S-s)\leq 2\epsilon^{-3/2}\cdot\mathrm{dist}(s,S-s)$.
@@ 10.3 Application: $1$-Means Range Query
Here, we use Lemma 10.1 to devise a data structure that supports estimating both the $1$-means objective and its solution on any given approximate ball. This data structure is obtained by plugging a dynamic coreset algorithm from [HenzingerK20] (for the entire set $X$) into Lemma 10.1; hence, for each query, we first obtain a small-sized coreset and then compute the objective and/or solution based on it.
@@
For any $0<\epsilon<1$, there exists a data structure that maintains a dynamic weighted dataset $X\subseteq[\Delta]^{d}$, and for every query point $x\in[\Delta]^{d}$ and radius $r\geq 0$, returns a real number $b_{x}\geq 0$, a center $c^{*}\in[\Delta]^{d}$, and two estimates $\widehat{\textsc{cost}}_{c^{*}}$ and $\widehat{\textsc{cost}}_{x}$, such that there exists a subset $B_{x}\subseteq X$ satisfying the following conditions.
@@
We first consider two extreme cases of the query radius $r$. In the case of a small radius $0\leq r<1$, we have $\textsc{Ball}(x,r)=\{x\}$ (recall that any two distinct points in the discrete space $\Delta]^{d}$ have distance at least 11), and we simply consider the point subset $B_{x}=\{x\}\cap X$. Either subcase – $B_{x}=\{x\}$ or $B_{x}=\emptyset$ – is trivial. In the case of a large radius $r>\sqrt{d}\Delta=\mathrm{diam}([\Delta]^{d})$, we have $\textsc{Ball}(x,r)\cap X=X$, and we simply consider the entire point set $B_{x}=X$. Thus, this extreme case $r>\sqrt{d}\Delta$ reduces to the boundary case $r=\sqrt{d}\Delta$.
@@
The following claim provides a data structure that (1) maintains the total weight of the dataset (a trivial task), and (2) constructs a coreset for the entire dataset, following the approach of [HenzingerK20].101010[HenzingerK20] provides a general $(1+\epsilon)$-coreset construction for $k$-means. In this paper, we only need a constant-factor coreset for $1$-means (e.g., $k=1$, $\epsilon=0.5$), so we state this special case.
@@
  • for every $j\in[t]$, the following holds with probability at least $\frac{1}{\mathrm{poly}(\Delta^{d})}.11$1111While Lemma 10.1 does not explicitly consider success probability, its analysis can be easily adapted to handle data structures that succeed with high probability.
@@ 11.1 Data Structure Description
In this part, we describe our data structure for maintaining an approximate assignment. We first present static and dynamic objects that are explicitly maintained by our data structure (Definition 11.1). Then, we show how these explicit objects define an implicit partition of the non-center points XSX-S (Claim 11.5). Built on top of this implicit partition, we define an implicit assignment $\sigma:X-S\to S$ (Claim 11.7).
@@
    $\big\{s\in S:\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq\tfrac{\rho_{i}}{\Gamma}\big\}\penalty 10000\ \subseteq\penalty 10000\ S(i,z)\penalty 10000\ \subseteq\penalty 10000\ \big\{s\in S:\mathrm{dist}(s,\varphi_{i}^{-1}(z))\leq 2\rho_{i}\big\}.$ (22)
@@
Initially, both $X$ and $S$ are empty $\emptyset$, so the initialization only needs to compute the $m+1$ hash functions $\{\varphi_{i}\}_{i\in[0,m]}$. By Lemma 9.2, every $\varphi_{i}$ can be constructed in $\tilde{O}(1)$ time with a success probability of $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$. Thus, the total initialization time is $(m+1)\cdot\tilde{O}(1)=\tilde{O}(1)$. Moreover, taking a union bound, the overall success probability is $1-\frac{m+1}{\mathrm{poly}(\Delta^{d})}=1-\frac{1}{\mathrm{poly}(\Delta^{d})}$. ∎
@@
Finally, we show how the explicit objects (Definition 11.1) implicitly define a partition of the non-center points XSX-S (Claim 11.5). This partition exhibits an approximate equidistant structure, where points in each part are roughly equally distant from the center set (Claim 11.6). We further define an implicit approximate assignment based on this partition (Claim 11.7).
@@
$\mathrm{dist}(s,x)$ $\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(s,\varphi_{i-1}^{-1}(z_{i-1}))+\mathrm{diam}(\varphi_{i-1}^{-1}(z_{i-1}))$ (Triangle inequality)
$\penalty 10000\ \leq\penalty 10000\ 2\rho_{i-1}+\rho_{i-1}$ (Item 2 of Definition 11.1)
$\penalty 10000\ =\penalty 10000\ \rho_{i}/\Gamma.$ (Item 1 of Definition 11.1)
Then, $\mathrm{dist}(s,\varphi_{i}^{-1}(z_{i}))\leq\mathrm{dist}(s,x)\leq\rho_{i}/\Gamma\implies s\in S(i,z_{i})=S(i,x)$ by Equations 22 and 23. The arbitrariness of $s\in S(i-1,x)$ implies that $S(i-1,x)\subseteq S(i,x)$. This finishes the proof. ∎
@@
$\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ is a partition of XSX-S.
@@
$\frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(x,S)\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(x,\sigma^{\prime}(i,z))\penalty 10000\ \leq\penalty 10000\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
@@
$\mathrm{dist}(\sigma^{\prime}(i,z),\varphi_{i}^{-1}(z))\penalty 10000\ \leq\penalty 10000\ 2\rho_{i}\penalty 10000\ =\penalty 10000\ (3\Gamma)^{i}.$
@@
$\mathrm{dist}(x,S)$ $\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(x,\sigma^{\prime}(i,z))$
$\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(\sigma^{\prime}(i,z),\varphi_{i}^{-1}(z))+\mathrm{diam}(\varphi_{i}^{-1}(z))$ ($\varphi_{i}(x)=z$)
$\penalty 10000\ \leq\penalty 10000\ (3\Gamma)^{i}+\frac{1}{2}\cdot(3\Gamma)^{i}$
$\penalty 10000\ =\penalty 10000\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
@@
$\min_{s\in S}\mathrm{dist}(s,\varphi_{i-1}^{-1}(z^{\prime}))\penalty 10000\ =\penalty 10000\ \frac{\rho_{i}}{\Gamma}\penalty 10000\ >\penalty 10000\ \frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}.$
@@
$\mathrm{dist}(x,S)\penalty 10000\ =\penalty 10000\ \min_{s\in S}\mathrm{dist}(x,s)\penalty 10000\ \geq\penalty 10000\ \min_{s\in S}\mathrm{dist}(s,\varphi_{i-1}^{-1}(z^{\prime}))\penalty 10000\ >\penalty 10000\ \frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}.$
@@
$\frac{1}{2\Gamma}\cdot(3\Gamma)^{i-1}\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(x,S)\penalty 10000\ \leq\penalty 10000\ \mathrm{dist}(x,\sigma^{\prime}(i,z))\penalty 10000\ \leq\penalty 10000\ \frac{3}{2}\cdot(3\Gamma)^{i}.$
@@
The following Claim 11.7 provides an implicit scheme for maintaining an approximate assignment, which follows directly from Claims 11.5 and 11.6.
@@ 11.2 Update Time Analysis
$\varphi_{i}(\textsc{Ball}(s,\tfrac{\rho_{i}}{\Gamma}))\penalty 10000\ \subseteq\penalty 10000\ \Phi_{i}\cap\varphi_{i}([\Delta]^{d})\penalty 10000\ \subseteq\penalty 10000\ \varphi_{i}(\textsc{Ball}(s,2\rho_{i})).$
@@
$z\penalty 10000\ =\penalty 10000\ \varphi_{i}(y)\penalty 10000\ \in\penalty 10000\ \varphi_{i}(\textsc{Ball}(x,(3\Gamma)^{i-1}))\penalty 10000\ \subseteq\penalty 10000\ \varphi_{i}(\textsc{Ball}(x,\tfrac{1}{\Gamma}\cdot(3\Gamma)^{i}))\penalty 10000\ \subseteq\penalty 10000\ \Phi_{x}.$
@@
The total update time for the data structure of Definition 11.1 follows by combining the results from Claims 11.9, 11.10, 11.11, 11.12 and 11.13. Among these, the most expensive one is maintaining the mappings $\{f(i,z)\}_{(i,z)\in H}$, which requires $\tilde{O}(\Lambda^{2})$ time. Therefore, the overall update time for the data structure is $\tilde{O}(\Lambda^{2})$. ∎